A defibrillator is used during a heart attack to restore the heart to its normal beating pattern. A defibrillator passes 18 A of current through the torso of a person in 1.7 ms. (a) How much charge moves during this time? (b) How many electrons pass through the wires connected to the patient?

Need the full answer for b. Not expessed as x10^
Answer #1

a)

Charge

Q=I*t =18*(1.7*10-3)

Q=0.0306 C

b)

N=Q/e =0,0306/(1.6*10-19)

N=1.9125*1017 electrons
Answer #2

The amount of charge moving in the time is,

                q = It

                   = (18 A)(1.7x10-3 s)

                  = 3.06x10-2 C

------------------------------------------------------------------

The number of electrons are,

          n = q/e

             = (3.06x10-2 C)/(1.6x10-19 C/electron)

             = 1.91x1017 electrons

             = 191250000000000000 electrons